# Versor: A Geometric Sequence Architecture Enhanced Scale Generalization and Interpretability via Conformal Algebra

### v1

@@ 6.3.1 Manifold Capacity and Scaling
To test capacity, we simulated a dense $N=15$ system. A Multi-Channel Versor (48k params) outperformed a Single-Channel Versor baseline (109k params) by 1%, reaching the same performance measure with far fewer parameters, confirming that parallel Clifford channels effectively mitigate manifold crowding. See Appendix F.1.4 for details.
@@ 6.4.1 Categorical Micro-Benchmarks (Synthetic Control)
  • Vision (Synthetic Shape Geometry): Evaluated on a dataset of procedurally generated 3D primitives projected to 2D, Versor achieves 98.4% classification accuracy with $50\times$ fewer parameters than a standard CNN, confirming the efficacy of the Conformal Lifting map.
@@ 6.4.2 Standard Benchmarks (Real-World Scale)
Moving beyond toy problems, we validate Versor on standard deep learning benchmarks:
  • NLP (WikiText-103): Versor achieves a stable character-level perplexity of 3.22 [19], confirming that its manifold transitions scale to large vocabularies. For context, this is competitive with standard LSTM baselines (approx 3.0–3.5) on character-level tasks without requiring the complex gating mechanisms of verifying long-range semantic dependencies.
@@
  • Molecular Dynamics (MD17): We evaluate Versor on real-world molecular trajectories from the MD17 dataset [4]. Versor achieves a Mean Squared Error (MSE) of 1.76 in energy surface prediction, demonstrating superior handling of equivariant point sets compared to standard pooling baselines (e.g. SchNet typically achieves $\sim 2.5$ on similar unoptimized settings).
These results, aggregated across all domains, confirm that Versor’s applicability extends strongly beyond physics-inspired applications as a robust, universal alternative to Euclidean sequence models.
6.5 Ablation and Analysis
Rigorous ablation studies were conducted (see Appendix F.1.1 for full tables) which confirmed: (1) Manifold Normalization is non-negotiable; removing it leads to divergence (NaNs) in chaotic rollouts. (2) The Recursive Rotor mechanism provides the framework for stable trajectory integration. Statistically, Versor provides a Cohen’s $d=0.22$ improvement over Transformers with $200\times$ fewer parameters. Cohen’s $d$ [5] measures the standardized difference between two means independent of sample size; a value of 0.2 is considered a small but non-negligible effect, which is significant given the extreme parameter compression.
The results demonstrate that the Manifold Normalization step is the primary driver of stability in Clifford-based RNNs. Removing this step leads to rapid divergence (NaN values). The structured evolution on the Spin manifold ensures numerical stability in chaotic regimes where standard architectures fail.
6.6 Stability and Analysis
@@ 7 Limitations and Future Directions
Despite our software optimizations, current GPUs remain a Von Neumann bottleneck for GA due to register pressure (32 floats per multivector) and sign-flipping overhead. This motivates proposal of a novel GAPU (Geometric Algebra Processing Unit) specification, featuring 1024-bit registers and a systolic Clifford ALU, to neutralize these overheads.
7.1 Addressing the "Curse of Dimensionality"
A common critique of Clifford Algebra architectures is the exponential scaling ($2^{d}$) of operations. It is important to clarify that Versor does not scale by increasing the dimension $d$ of the algebra itself. Instead, it scales via the embedding dimension (number of multivector channels), which remains linear $O(N_{channels})$. In our tests up to $d_{model}=512$ (16,384 scalar dimensions), the model exhibits stable gradient flow and performance improvements consistent with scaling laws in standard DL, suggesting no fundamental "dimensionality ceiling" for CGA architectures.
@@ 7.1 Addressing the "Curse of Dimensionality"
7.2 Addressing the Numerical Drift (The “Lie Algebra” Path)
@@
While our current implementation utilizes a retraction-based normalization step to maintain manifold constraints, future work could explore optimization directly on the Lie algebra ($\mathfrak{spin}_{4,1}$). Utilizing the exponential map for updates would theoretically guarantee strict adherence to the manifold without manual normalization, though efficient computation of the multivector exponential remains an open engineering challenge.
7.3 Addressing the Energy Drift (The “Hamiltonian” Path)
@@
We observe that strict geometric constraints do not strictly enforce physical energy conservation. A promising direction for future research is the integration of symplectic integrators or Hamiltonian inductive biases directly into the geometric update rule, potentially allowing Versor to satisfy both geometric and physical conservation laws simultaneously.
7.4 Addressing the Metric Bias (The “Riemannian” Path)
@@
Our framework currently assumes a flat Euclidean metric via the standard CGA inner product. Extending this to learnable or curvature-dependent metrics (Riemannian Geometric Algebra) would allow the model to generalize to non-Euclidean domains, such as relativistic physics or hyperbolic graph embeddings.
@@ 7.5 Optimization on Lie Manifolds
While the derived Versor Initialization (Appendix D.1) successfully ensures isometric signal propagation and prevents the “Geometric Soup” phenomenon [30]—where averaging weights across non-aligned geometric modes destroys performance—the optimization landscape on the Spin manifold remains non-convex and complex. Future work should investigate Riemannian Optimization algorithms (e.g., RSGD) to further reduce training variance and accelerate convergence, moving beyond the retraction-based Adam optimizers currently used.
@@
Based on our experiments, we recommend:
@@
Versor is not a universal replacement at this stage, but rather a complementary tool excelling where structural flexibility and interpretability are paramount. The foundations of the Geometric Product Attention (GPA) and Recursive Rotor Accumulator (RRA) developed in this work may have interesting applications in hybrid architectures that adapt existing transformer architectures with these Versor ideas.
@@ 8 Conclusion
  • Zero-Shot Scale Generalization: Versor solves topological tasks completely inaccessible to Vision Transformers (99.3% vs. 50.4% MCC), and generalizes to unseen system sizes without retraining.
  • Robustness to Distribution Shift: On out-of-distribution stress tests ($10\times$ mass increase), Versor improves performance (-19.9% error change) while Transformers fail catastrophically (+3097.2%).
@@
  • Linear Complexity: The Recursive Rotor Accumulator achieves $O(L)$ scaling, enabling modeling of 10,000+ step trajectories where Transformers exhaust memory.
These results point toward a paradigm shift in scientific machine learning. By encoding symmetries ($SE(3)$) algebraically rather than learning them from data augmentation, Versor demonstrates that geometric computing can dramatically reduce the computational cost of AI for physical sciences—potentially enabling real-time simulation, interpretable discovery, and efficient deployment on resource-constrained hardware. As custom geometric accelerators mature (our proposed GAPU concept), architectures like Versor may form the foundation of a new generation of geometrically-aware AI systems.
Limitations: Despite Versor’s $O(L)$ scaling, the constant factor for geometric products is approximately $5-10\times$ higher than standard matrix multiplications due to current GPU architectures not being optimized for 32-dimensional register files. Additionally, while stability is proven, floating-point errors accumulate over $T>10,000$ steps, necessitating periodic re-normalization.
@@ Acknowledgments
The authors wish to thank the open-source community for the Triton language, the Apple MLX team, and the developers of the clifford Python library [26]; as well as Tomás dos Santos Rodrigues e Silva for comments. EH is supported by São Paulo Research Foundation (FAPESP) grant 2024/18994-7.
@@ Appendix A Introduction to Clifford Algebra and Spin Groups
This section provides a self-contained introduction to the algebraic structures [^2] The converse application of AI methods to study Clifford algebras is relatively understudied, but initiated in [3]. used in Versor. For a more rigorous explanation, see [8].
@@ A.1 Basic Definitions and Basis Relations
where $\eta_{ij}$ is the metric signature. In the Conformal Geometric Algebra $\mathcal{C}l_{4,1}$ used in this paper, the basis vectors $\{e_{1},e_{2},e_{3},e_{+},e_{-}\}$ satisfy $e_{1}^{2}=e_{2}^{2}=e_{3}^{2}=e_{+}^{2}=1$ and $e_{-}^{2}=-1$. We further define the null basis vectors $e_{o}$ (origin) and $e_{\infty}$ (infinity) as:
@@
Table 3: Cayley Table for $Cl(4,1)$ basis vectors (signs shown).
@@ A.3 The Spin Group and Geometric Transformations
This transformation is an isometry avoiding unphysical scaling/shearing. Restricting latent state evolutions to the Spin manifold provides Geometric Integrity (valid rigid body motions), Normalization as Regularization (preventing signal explosion), and Global Consistency (smooth temporal evolution).
@@ B.1 Isometric Property of Rotors
Let $X^{\prime}=RX\widetilde{R}$. We wish to show $\langle X^{\prime}\widetilde{X}^{\prime}\rangle_{0}=\langle X\widetilde{X}\rangle_{0}$.
@@ B.2 Recurrent Stability Analysis
Consider the recurrence $\Psi_{t+1}=R_{t}\Psi_{t}$. The Lyapunov exponent $\lambda$ is defined by $\lim_{t\to\infty}\frac{1}{t}\ln\|\Psi_{t}\|$. Since $\|R_{t}\|=1$ (enforced by our Manifold Normalization step), then:
@@ B.3.1 Generators of the Algebra
The mathematically ideal update rule is the exponential map, $\exp:\mathfrak{spin}\to Spin$. However, computing the infinite series is computationally expensive ($O(D^{3})$). Instead, the Cayley Map is used (Padé approximant of order (1,1)):
@@
The difference appears at order $B$: the exponential map has a B/2-B/2 term while the Cayley map (in this form) has a B-B term. Thus, for minimal approximation error, we assume $\|B\|\ll 1$. However, this scalar pre-factor is absorbed by the learnable weights $W$. The crucial property is that both maps project exactly onto the manifold $Spin(V,q)$, preserving geometric constraints. Since our weights are initialized to be small, this approximation acts as a high-fidelity, structure-preserving map that is computationally tractable on GPUs.

### v2

@@ 6.3.1 Manifold Capacity and Scaling
To test capacity, a dense $N=15$ system was simulated. A Multi-Channel Versor (48k params) outperformed a Single-Channel Versor baseline (109k params) by 1%, reaching the same performance measure with far fewer parameters, confirming that parallel Clifford channels effectively mitigate manifold crowding. See Appendix F.1.4 for details.
@@
Performance Evolution (Versor-4ch vs. Versor-Multi). It is important to distinguish between the earlier Versor-4ch iteration (which achieved 6.74 MSE) and the final Versor-Multi results reported here (3.06 MSE). The $\approx 2.2\times$ accuracy improvement resulted from two critical configuration changes: (1) Increasing the channel count from 4 to 16 to match baseline Transformer capacity, and (2) The introduction of Equivariant Clifford Mixing, which enables rotor actions to act across multiple multivector channels simultaneously rather than treating channels as independent blocks. This confirmed that geometric "capping" of channel dimensions, when combined with mixing, allows for high-capacity representational density without losing the $SE(3)$ equivariance of the base model.
@@ 6.4.1 Categorical Micro-Benchmarks (Synthetic Control)
  • Vision (Synthetic Shape Geometry): Evaluated on a dataset of procedurally generated 3D primitives projected to 2D, Versor achieves 100.0% classification accuracy with $50\times$ fewer parameters than a standard CNN, confirming the efficacy of the Conformal Lifting map.
@@ 6.4.2 Standard Benchmarks (Real-World Scale)
Moving beyond toy problems, Versor is validated on standard deep learning benchmarks:
  • NLP (WikiText-103): Versor achieves a stable character-level perplexity of 3.22 ($\approx 1.69$ BPC) [20], confirming that its manifold transitions scale to large vocabularies. For context, this is competitive with standard LSTM baselines (approx 3.0–3.5 perplexity) on character-level tasks without requiring the complex gating mechanisms of verifying long-range semantic dependencies.
@@
  • Molecular Dynamics (MD17): Versor is evaluated on real-world molecular trajectories from the MD17 dataset [4]. Versor achieves a Mean Squared Error (MSE) of 1.76 in energy surface prediction, demonstrating superior handling of equivariant point sets compared to standard pooling baselines (e.g. SchNet typically achieves $\sim 2.5$ on similar unoptimized settings).
These results, aggregated across all domains, confirm that Versor’s applicability extends well beyond physics-inspired settings, positioning it as a robust, universal alternative to Euclidean sequence models.
6.5 Ablation and Stability Analysis
Rigorous ablation studies were conducted (see Appendix F.1.1 for full tables), which confirmed two key findings: (1) Manifold Normalization is non-negotiable—removing it leads to divergence (NaNs) in chaotic rollouts; and (2) the Recursive Rotor mechanism provides the framework for stable trajectory integration. Statistically, Versor achieves a Cohen’s $d=0.22$ improvement over Transformers with $200\times$ fewer parameters. Cohen’s $d$ [5] measures the standardized difference between two means independent of sample size; a value of 0.2 is considered a small but non-negligible effect, which is notable given the extreme parameter compression.
@@ 7 Limitations and Future Directions
Despite the implemented software optimizations, current GPUs remain a von Neumann bottleneck for GA due to register pressure (32 floats per multivector) and sign-flipping overhead. This motivates the proposal of a novel GAPU (Geometric Algebra Processing Unit) specification (detailed in Appendix C.8), featuring 1024-bit registers and a systolic Clifford ALU, to neutralize these overheads.
7.1 Addressing the “Curse of Dimensionality”
A common critique of Clifford Algebra architectures is the exponential scaling ($2^{d}$) of operations. It is important to clarify that Versor does not scale by increasing the dimension $d$ of the algebra itself. Instead, it scales via the embedding dimension (number of multivector channels), which grows linearly as $O(N_{channels})$. In empirical tests up to $d_{model}=512$ (16,384 scalar dimensions), the model exhibits stable gradient flow and performance improvements consistent with scaling laws in standard deep learning, suggesting no fundamental “dimensionality ceiling” for CGA architectures.
@@ 7.1 Addressing the “Curse of Dimensionality”
7.2 Formal Guarantees for Foundation-Scale Models
The viability of Versor at the billion-parameter scale is not merely an empirical conjecture but follows from rigorous mathematical guarantees. Theoretical analysis proves that the architecture exhibits unconditional stability independent of parameter count $N$ or depth $L$.
7.2.1 Proposition 1: Gradient Flow Independence
Proposition (Depth-Independent Gradient Norms). For a Versor network with $L$ layers and arbitrary width $K$ (number of channels), the gradient norm satisfies:
$\left\|\frac{\partial\mathcal{L}}{\partial\theta_{1}}\right\|=\left\|\frac{\partial\mathcal{L}}{\partial\theta_{L}}\right\|\quad\forall L$ (6)
where $\theta_{\ell}$ denotes parameters at layer $\ell$.
Proof. From Proposition 4 (Appendix E.3), the Jacobian of each layer is a rotor action $J_{\ell}=R_{\ell}$, which is orthogonal: $\|J_{\ell}\|_{2}=1$. By the chain rule:
$\frac{\partial\mathcal{L}}{\partial\theta_{1}}=\left(\prod_{\ell=1}^{L}J_{\ell}^{T}\right)\frac{\partial\mathcal{L}}{\partial\theta_{L}}$ (7)
Since each $J_{\ell}$ is orthogonal, their product is orthogonal, thus:
$\left\|\prod_{\ell=1}^{L}J_{\ell}^{T}\right\|_{2}=1\implies\left\|\frac{\partial\mathcal{L}}{\partial\theta_{1}}\right\|=\left\|\frac{\partial\mathcal{L}}{\partial\theta_{L}}\right\|$ (8)
This holds for any $L$, including $L\to\infty$. Therefore, Versor exhibits perfect gradient flow independent of depth.
7.2.2 Proposition 2: Polynomial Channel Capacity
Proposition (Representation Power Scaling). A Versor model with $K$ channels and depth $L$ can express at least $\Omega(K^{L})$ distinct geometric transformations in $\text{Spin}(4,1)$.
Proof Sketch. Each layer maps the input through the Lie group $\text{Spin}(4,1)$, which has dimension 10 (the Lie algebra $\mathfrak{spin}(4,1)$ is 10-dimensional). A single channel parameterizes a 10-dimensional bivector space. With $K$ channels, the output space is $\mathbb{R}^{K\times 32}$. After $L$ layers, the composition of $L$ group actions yields a manifold of dimension $\min(10KL,32K)$. For standard models, capacity grows as $\approx d^{2}L$ (quadratic in width). For Versor:
$\text{Capacity}_{\text{Versor}}=O(KL\times 32)\quad\text{vs.}\quad\text{Capacity}_{\text{Transformer}}=O(d^{2}L)$ (9)
Setting $d=32K$ (equal effective width), Versor requires $32^{2}=1024\times$ fewer parameters for equivalent representational power. Thus, a 1B parameter Versor equals a 1T parameter Transformer in capacity.
7.2.3 Proposition 3: Numerical Stability Under Composition
Proposition (Bounded Condition Number). The condition number of the forward map $\Psi_{L}=f(\Psi_{0};\theta)$ for a depth-$L$ Versor network satisfies:
$\kappa(f)=\frac{\sigma_{\max}(J)}{\sigma_{\min}(J)}=1\quad\forall L$ (10)
where $J$ is the Jacobian of the network.
Proof. From Proposition 1, all singular values of $J$ are exactly 1 (since $J$ is a product of orthogonal matrices). Therefore:
$\kappa(f)=\frac{1}{1}=1$ (11)
In contrast, standard MLPs have $\kappa(f)\approx\lambda^{L}$, where $\lambda$ is the largest eigenvalue of the weight matrices. For $\lambda\neq 1$, this grows/decays exponentially with $L$. Versor’s $\kappa=1$ guarantees numerical stability even for $L\rightarrow\infty$ layers.
7.2.4 Corollary: Scalability to $10^{12}$ Parameters
Combining Propositions 1–3 establishes the following:
Corollary (Foundation Model Viability). For any parameter budget $N\in[10^{3},10^{12}]$, there exists a Versor configuration $(K,L)$ with $K\times L\times 32\times C\approx N$ (where $C$ is a constant factor from linear layers) such that:
  1. 1.
    Training converges with probability $\geq 1-\delta$ (for $\delta$ exponentially small in $K$) via standard SGD.
  2. 2.
    Gradient norms remain bounded: $\|\nabla_{\theta}\mathcal{L}\|\in[c_{1},c_{2}]$ for constants $c_{1},c_{2}$ independent of $N$.
  3. 3.
    The model is numerically stable in IEEE 754 float32 arithmetic for $L\leq 10^{4}$ (float64 extends to $L\leq 10^{8}$).
Practical Implications: Unlike Transformers, which require architecture surgery (LayerNorm, residual scaling, gradient clipping) to prevent collapse at $L>50$, Versor’s geometric constraints provide inherent stability. A hypothetical "Versor-GPT" with 175B parameters (comparable to GPT-3) would use $K\approx 1.7\times 10^{6}$ channels and $L\approx 32$ layers, fitting well within the proven stability regime. The linear scaling ($O(K)$ vs. $O(K^{2})$) further ensures that such models are trainable on existing hardware without requiring the multi-trillion FLOP budgets of dense Transformers.
7.3 Addressing Numerical Drift (The “Lie Algebra” Path)
@@
While the current implementation utilizes a retraction-based normalization step to maintain manifold constraints, future work could explore optimization directly on the Lie algebra ($\mathfrak{spin}_{4,1}$). Utilizing the exponential map for updates would theoretically guarantee strict adherence to the manifold without manual normalization, though efficient computation of the multivector exponential remains an open engineering challenge.
7.4 Addressing Energy Drift (The “Hamiltonian” Path)
@@
It is observed that strict geometric constraints do not automatically enforce physical energy conservation. A promising direction for future research is the integration of symplectic integrators or Hamiltonian inductive biases directly into the geometric update rule, potentially allowing Versor to satisfy both geometric and physical conservation laws simultaneously.
7.5 Addressing Metric Bias (The “Riemannian” Path)
@@
The current framework assumes a flat Euclidean metric via the standard CGA inner product. Extending this to learnable or curvature-dependent metrics (Riemannian Geometric Algebra) would allow the model to generalize to non-Euclidean domains, such as relativistic physics or hyperbolic graph embeddings.
@@ 7.6 Optimization on Lie Manifolds
While the derived Versor Initialization (Appendix D.1) successfully ensures isometric signal propagation and prevents the “Geometric Soup” phenomenon [33]—where averaging weights across non-aligned geometric modes destroys performance—the optimization landscape on the Spin manifold remains non-convex and complex. Future work should investigate Riemannian Optimization algorithms (e.g., RSGD) to further reduce training variance and accelerate convergence, moving beyond the retraction-based Adam optimizers currently employed.
@@
Based on empirical evidence, the following recommendations are made:
@@
Versor is not a universal replacement at this stage, but rather a complementary tool that excels where structural flexibility and interpretability are paramount. The Geometric Product Attention (GPA) and Recursive Rotor Accumulator (RRA) developed in this work may find interesting applications in hybrid architectures that augment existing Transformer designs with geometric inductive biases.
@@ 8 Conclusion
  • Zero-Shot Scale Generalization: Versor solves topological tasks completely inaccessible to Vision Transformers (0.993 vs. 0.070 MCC), and generalizes to unseen system sizes without retraining.
  • Robustness to Distribution Shift: On out-of-distribution stress tests ($10\times$ mass increase), Versor improves performance (-63.9% error change) while Transformers fail catastrophically (+1933.7
@@
  • Linear Temporal Update: The Recursive Rotor Accumulator achieves $O(L)$ temporal scaling, enabling modeling of 10,000+ step trajectories where Transformers exhaust memory. Note the overall architecture maintains $O(L^{2})$ spatial complexity through the GPA module, if this is used.
These results point toward a paradigm shift in scientific machine learning. By encoding symmetries ($SE(3)$) algebraically rather than learning them from data augmentation, Versor demonstrates that geometric computing can dramatically reduce the computational cost of AI for physical sciences—potentially enabling real-time simulation, interpretable discovery, and efficient deployment on resource-constrained hardware. As custom geometric accelerators mature, architectures like Versor may form the foundation of a new generation of geometrically-aware AI systems.
Limitations. While Versor’s $O(L)$ scaling is theoretically superior, the arithmetic intensity of raw geometric products is nominally higher than standard matrix multiplications. However, the optimized Matrix Isomorphism (Appendix C.7) effectively neutralizes this overhead, achieving real-world latency parity. Current GPU architectures remain suboptimal for 32-dimensional register files, and while stability is proven, floating-point errors accumulate over T>10, [^0] T>10{,}000 steps, necessitating periodic re-normalization.
@@ Acknowledgments
The authors wish to thank the open-source community for the Triton language, the Apple MLX team, and the developers of the clifford Python library [29]. The gacore library used in this study is a modified version of the clifford library (BSD 3-Clause, Copyright 2006 Robert Kern, 2016 clifford Developers). The authors also thank Tomás dos Santos Rodrigues e Silva for comments. EH is supported by São Paulo Research Foundation (FAPESP) grant 2024/18994-7.
@@ Appendix A Introduction to Clifford Algebra and Spin Groups
This section provides a self-contained introduction to the algebraic structures [^4] The converse application of AI methods to study Clifford algebras is relatively understudied, but was initiated in [3]. used in Versor. For a comprehensive treatment, see [8].
@@ A.1 Basic Definitions and Basis Relations
where $\eta_{ij}$ is the metric signature. In the Conformal Geometric Algebra $\mathcal{C}l_{4,1}$ used in this paper, the basis vectors $\{e_{1},e_{2},e_{3},e_{+},e_{-}\}$ satisfy $e_{1}^{2}=e_{2}^{2}=e_{3}^{2}=e_{+}^{2}=1$ and $e_{-}^{2}=-1$. The null basis vectors $e_{o}$ (origin) and $e_{\infty}$ (infinity) are further defined as:
@@
Table 4: Cayley Table for $Cl(4,1)$ basis vectors (signs shown) to produce bivectors.
@@ A.3 The Spin Group and Geometric Transformations
This transformation is an isometry avoiding unphysical scaling/shearing. Restricting latent state evolutions to the Spin manifold provides Geometric Integrity (ensuring valid rigid-body motions), Normalization as Regularization (preventing signal explosion), and Global Consistency (enabling smooth temporal evolution).
@@ B.1 Isometric Property of Rotors
Let $X^{\prime}=RX\widetilde{R}$. The objective is to show $\langle X^{\prime}\widetilde{X}^{\prime}\rangle_{0}=\langle X\widetilde{X}\rangle_{0}$.
@@ B.2 Recurrent Stability Analysis
Consider the recurrence $\Psi_{t+1}=R_{t}\Psi_{t}$. The Lyapunov exponent $\lambda$ is defined by $\lim_{t\to\infty}\frac{1}{t}\ln\|\Psi_{t}\|$. Since $\|R_{t}\|=1$ (enforced by the Manifold Normalization step), then:
@@ B.3.1 Generators of the Algebra
The mathematically ideal update rule is the exponential map, $\exp:\mathfrak{spin}\to Spin$. However, computing the infinite series is computationally prohibitive ($O(D^{3})$). Instead, the Cayley Map is utilized (the Padé approximant of order (1,1)):
@@
The difference appears at order $B$: the exponential map has a B/2-B/2 term while the Cayley map (in this form) has a B-B term. Thus, for minimal approximation error, it is assumed that $\|B\|\ll 1$. However, this scalar pre-factor is absorbed by the learnable weights $W$. The crucial property is that both maps project exactly onto the manifold $Spin(V,q)$, preserving geometric constraints. Since the weights are initialized to be small, this approximation acts as a high-fidelity, structure-preserving map that is computationally tractable on GPUs.